Compute the number of vinyl tiles, measuring 6 in. each side, needed to tile a kitchen measuring 20 ft. by 16 ft.
How many tiles?

Answer by edjones(8007)

20*12=240"
16*12=192"
192*240=46080"
46080/36=1280 tiles
.
Ed